OpenStudy (aaronandyson):

sec^2(A) + cosec^2(A) = sec^2(A).cosec^2(A) @michele_liano

OpenStudy (aaronandyson):

@Michele_Laino

OpenStudy (michele_laino):

we have to apply these identities: \[\sec A = \frac{1}{{\cos A}},\quad \csc A = \frac{1}{{\sin A}}\]

OpenStudy (aaronandyson):

?

imqwerty (imqwerty):

have u tried to solve the expression?? try converting the terms into sin nd cos form nd then take the LCM nd solve

OpenStudy (michele_laino):

left side becomes: \[\sec A + {\left( {\csc A} \right)^2} = \frac{1}{{\cos A}} + {\left( {\frac{1}{{\sin A}}} \right)^2} = ...\] please continue

OpenStudy (aaronandyson):

i'm confused

OpenStudy (michele_laino):

I think that there is a typo into your original expression, please check

OpenStudy (aaronandyson):

fixed

OpenStudy (michele_laino):

ok! so left side becomes: \[{\left( {\sec A} \right)^2} + {\left( {\csc A} \right)^2} = {\left( {\frac{1}{{\cos A}}} \right)^2} + {\left( {\frac{1}{{\sin A}}} \right)^2} = ...\]

imqwerty (imqwerty):

ok m calling sinA = x nd cosA=y 1/y^2 +1/x^2 take LCM we get - (x^2 + y^2)/(xy)^2 we know that sin^2A+cos^2A = 1 therefore the numerator = 1 nd we r left with 1/(xy)^2 =sec^A cosec^A

OpenStudy (aaronandyson):

sin^2(A)+cos^2(A)/cos(A)sin(A) = RHS

imqwerty (imqwerty):

yes

OpenStudy (aaronandyson):

1/sin(A)cos(A) = sec(A)cosec(A) = RHS

OpenStudy (michele_laino):

more precisely, it is: sin^2(A)+cos^2(A)/cos(A)^2 sin(A)^2 = RHS

OpenStudy (aaronandyson):

ok!thanks...

OpenStudy (aaronandyson):

i nedd more help *need